You can buy commercial paper of a major US corporation for $495,000. The paper has a face value of $500,000 and is 45 days from maturity. Calculate the discount yield and bond equivalent yield on the commercial paper.
Answer #1

Discount Yield

D= (500000-495000)/500000*(360/45) = 0.08 i.e 8%

Bond Yield

B = (500000-495000)/495000*(365/45) = 0.08193 i.e 8.19%